OpenStudy (anonymous):

Solve the system by elimination: 4x + 8y = 20 -4x + 2y = -30

jimthompson5910 (jim_thompson5910):

Add the equations to get 4x + 8y = 20 -4x + 2y = -30 -------------- 0x+10y = -10 So we then get 10y = -10

jimthompson5910 (jim_thompson5910):

Solve 10y = -10 for y Then use this solution to find the solution for x.